# Large Language Model for Qualitative Research - A Systematic Mapping Study

### v1

@@ Abstract
The presence of Artificial Intelligence (AI) has been intensifying in supporting various human activities. Generative Artificial Intelligence (Generative AI) creates new data from learned patterns, facilitating the automation of complex processes. A promising application of this technology is Large Language Models (LLMs), which have proven effective in analyzing large volumes of text, enabling the extraction and understanding of information [1][2][3].
LLMs are specialized Generative AI models designed for text processing and generation. The term ’large’ refers to the massive number of parameters and the extensive textual data used in their training, allowing these models to identify complex patterns in natural language. Based on these patterns, LLMs can perform various tasks, such as machine translation, summarization, text editing, chatbot assistance, and even code generation for software development. These capabilities arise from LLMs ability to understand context and adjust their responses to different demands, making them powerful tools for automating tasks that previously required human intervention [4][5].
One of the prominent fields for applying LLMs is qualitative analysis. Qualitative analysis is a research approach aimed at exploring and interpreting collected data to identify, develop, and categorize concepts based on their properties and dimensions. This process involves segmenting the data into smaller parts, continuously comparing them, and assigning labels that accurately synthesize observed events. Through this method, raw data is transformed into abstract concepts, allowing the formation of broader categories that enable a deeper understanding of the phenomena being studied [6].
In software engineering, qualitative analysis using LLMs becomes a powerful tool by providing an analytical framework to organize large volumes of unstructured textual data. This facilitates the identification of underlying patterns and the development of categories that help in understanding complex problems. By avoiding biases and premature interpretations, researchers can uncover valuable insights that aid in the formulation of new theories and the enhancement of analytical processes, especially in contexts where the complexity of data requires a thorough and systematic approach to ensure effective and applicable results [6].
LLMs can significantly reduce the effort and time required for qualitative analysis of large data volumes by automating steps that traditionally demand exhaustive human analysis.
@@
This paper is structured as follows: Section II provides a brief background on qualitative analysis and LLMs. Section III describes the systematic mapping protocol and results. Section IV presents a summary of contributions and research opportunities. Section V discusses the threats to validity. Finally, Section VI presents the main conclusions.
@@ II-B Capabilities of LLMs in Overcoming Limitations
LLMs offer transformative solutions to many challenges faced by traditional qualitative research methods. LLMs, such as GPT-4, LLaMA, and ChatGPT, are generative AI systems trained on vast datasets, enabling them to understand and generate human-like text [4], [5]. By leveraging billions of parameters, these models excel in identifying patterns, performing thematic analysis, and automating coding processes.
@@
Reducing Subjectivity: While human coding is prone to inconsistency, LLMs provide consistent and reproducible outputs, especially when fine-tuned or used with standardized prompts [14]. This minimizes bias and ensures greater reliability in qualitative findings.
@@ II-C Illustrative Examples
Healthcare Case Study: In a study analyzing patient feedback, an open-source LLM was fine-tuned to perform inductive thematic analysis, identifying critical concerns like delays in care and communication breakdowns [17]. This process significantly reduced the time required for analysis while maintaining accuracy.
@@ II-D Motivation for This Study
Despite their potential, LLMs face limitations that must be addressed for optimal application in qualitative research. For example, they are highly dependent on well-structured prompts, which, if poorly designed, can lead to inaccurate or incomplete outputs [19], [17]. Additionally, LLMs can generate ”hallucinations”—fabricated responses that lack basis in the data [13]. These challenges highlight the need for systematic studies to explore best practices, identify gaps, and refine methodologies for integrating LLMs into qualitative research.
@@ III-B Search Strategy
A generic search string was developed using the keywords ”LLM” and ”Qualitative Analysis”. These keywords were connected using the logical operator AND, while their synonyms were linked with the OR operator. The terms in the search string were selected to ensure broader coverage of relevant studies. The string was tested in various configurations on the Scopus database, and after a calibration process, the final version defined was:
(”LLM” OR ”LLMs” OR ”Large Language Model” OR ”Large Language Models”) AND (”Qualitative Analysis” OR ”Qualitative Research” OR ”Grounded Theory”)
In addition to Scopus222http://www.scopus.com, searches were also conducted in the ACM Digital Library333https://dl.acm.org/, IEEExplore444https://ieeexplore.ieee.org/, Web of Science555http://www.isiknowledge.com, and SBC Open Lib databases666https://sol.sbc.org.br/busca/. Considering the recent nature of the topic and the possibility of studies still awaiting peer review, the Arxiv777https://arxiv.org/ database was also consulted. The searches were carried out during the months of September and October 2024, using filters applied to titles, abstracts, and keywords. No restrictions were placed on the starting publication year. In total, 354 studies were retrieved.
@@ III-E Conducting
Subsequently, a refinement process was conducted to select the final set of studies. First, duplicate studies were removed, resulting in 253 studies. Next, inclusion and exclusion criteria were applied based on the reading of titles and abstracts, reducing the total to 34 studies. In the following stage, only studies available for download were included, which further reduced the count to 21.
If a study did not fully meet the response for criterion DE1 (i.e., the response was ”no” or ”partially”), the remaining criteria were not evaluated, and the study was excluded. Additionally, if a study did not meet criterion DE11 (”no”), it was excluded for not describing the prompt engineering used, which is essential to ensure reproducibility.
At the end of this process, 7 studies remained, for which comprehensive data extraction was conducted. Table III presents all the studies that were included in this systematic mapping.
@@ III-F Results
During the analysis of the studies, it was found that 5 out of the 7 included works (approximately 71.4%) were published in 2024, while the remaining 2 were published in 2023. These data highlight the recent nature of the topic related to the evaluation of the use of LLMs for qualitative analysis, while also reflecting the growing interest of the scientific community and the increase in the number of publications on the subject.
@@
The studies exploring the use of LLMs as a tool to support qualitative data analysis are concentrated in various fields, including healthcare [14] [17], education [13][15][19], cultural studies [18], and analysis of technological applications [16].
Regarding the primary objective of the studies applying LLMs in qualitative analysis, three of them aimed to evaluate the capability and potential of LLMs for automating this type of analysis [13][15][19]. In studies [14][18], the main focus was on automating the qualitative data analysis process. Meanwhile, the works by [16][17] aimed to compare the results of qualitative analysis assisted by LLMs with those obtained through manual analysis.
@@
Regarding the LLM models tested in the studies, the following were used: ChatGPT [14] [16] [19], LLaMA-2 [17], ATLAS.TI [13], BERT [15], and the Sabiá-2 medium model [18]. Approximately 57.1% of the studies (4 out of 7) reported making adjustments or adaptations to the LLM to carry out the analyses [13] [14] [15] [17].
As for the data sources used in the included studies, most employed data from interviews [13] [17] [19], followed by those that analyzed documents [14] [15], social media (user reviews of applications) [16], and song lyrics [18].
@@
All included studies employed techniques and/or methodologies to support qualitative data analysis, also detailing the prompt engineering used. Two studies applied theme extraction combined with content analysis [13] [16]. In other studies, the techniques of open coding and theme extraction were used in conjunction with content analysis [17] [19]. Two studies combined categorization and theme extraction techniques; one integrated the Grounded Theory method [14], and the other used content analysis [15]. In study [18], three techniques were tested: categorization, theme extraction, and topic extraction.
Regarding how LLMs were used to conduct the qualitative analysis process, the approaches varied. Three studies utilized prompt-based instruction for the LLM [16] [17] [19]. In study [18], the authors combined prompt-based instruction with automated extraction using the BERT model, while study [15] exclusively conducted automated extraction using the BERT model. The remaining studies employed automated extraction using ATLAS.TI, associated with a specific technique [13], or conducted model training [14].
@@
All seven included studies conducted comparisons between qualitative analysis performed with the support of LLMs and traditional analysis methods. In some cases, such as in studies [13] [14] [15] [17] [19], the use of LLMs yielded results classified as equivalent to traditional methods. In study [18], the performance of the LLM surpassed that of traditional methods. Conversely, in study [16], the final result was considered inferior to that of traditional analysis, particularly when compared to human analysis.
The metrics used to evaluate performance included comparisons between LLM-assisted analysis and human analysis [13] [16] [17] [19], comparisons with human analysis and accuracy [15], and comparisons involving human analysis, accuracy, F1 Score, and precision [14] [18].
@@
All the included studies reported specific limitations regarding the use of LLMs in performing qualitative analyses. One limitation mentioned was the dependence on well-structured prompts to achieve accurate and reliable results [17] [19]. Another limitation relates to the tendency of LLMs to generate ”hallucinations,” meaning responses that lack a clear justification in the data [13] [19]. Additionally, the risk of inherent biases in the models was noted, especially when dealing with sensitive information [17]. Other limitations included difficulties in assigning topics to subjective expressions [18], limitations in context sensitivity and emotional nuances [16], and restrictions in interpreting complex meanings [19]. According to [14], there was also a tendency toward over-segmentation of responses and excessive classification into specific categories.
Additional reported limitations involve the AI’s difficulty in performing broader and more detailed categorization, which results in the loss of specific details considered important by human researchers [13], as well as inconsistent behaviors in zero-shot classifications [18]. Another highlighted issue is the tendency of LLMs to create an ”echo chamber” effect by reproducing pre-trained patterns [16].
Regarding recommendations for the use of LLMs in qualitative analysis, studies [13] [17] [19] recommend that LLMs be used as an aid, rather than as a replacement for human analysts, suggesting their use, for example, in categorization stages, with researchers responsible for the final interpretation of the data. Additionally, these studies emphasize the importance of human collaboration and the use of LLMs to enhance accuracy and increase the validity of results [16] [19].
In terms of suggestions for future research directions, the studies point to the need for improvements in prompt engineering techniques, recommending the exploration of more complex architectures to enhance the textual understanding of LLMs [15] [16] [18] [19]. There was also a recommendation to increase the volume of training data to improve accuracy [15], as well as to fine-tune pre-trained models for better topic assignments [18]. Other suggested future directions include enhancing LLM capabilities to capture semantic and subjective nuances [13], continuously improving open-source models to reduce costs and improve accessibility, and developing more user-friendly interfaces to facilitate their usage [17].
@@ IV Summary of Contributions and Research Opportunities
This section summarizes the main results and contributions obtained in this study, as well as the research opportunities identified through this mapping. The key contributions of this paper include:
Mapping of the field: This study provides an overview of the use of LLMs for qualitative analysis, addressing various dimensions of the area, such as: (i) the contexts in which LLMs have been applied to support qualitative analysis, (ii) the LLM models used and the data sources employed, (iii) the techniques and methodologies adopted in conducting qualitative analysis with the help of LLMs, (iv) the metrics used to evaluate effectiveness, and (v) the limitations and research opportunities identified in the current state of the art.
This mapping focused on studies that reported the use of LLMs for qualitative analysis and detailed the prompt engineering employed, which is an essential requirement to ensure replicability. The results highlight the recent nature of exploration in this field, reflected by the relatively low number of included studies, most of which were published recently in 2023 and 2024.
However, the findings allowed for the identification of several research opportunities. While LLMs demonstrate performance close to or even superior to traditional methods in some cases, there is a clear need for improvements. One aspect that requires further exploration is the limitations associated with the reliance on well-structured prompts. In this regard, it would be relevant to explore approaches that enhance LLMs’ textual understanding, enabling greater flexibility in qualitative analyses.
Additionally, there are opportunities to investigate new architectures or techniques that could improve the capture of semantic and subjective nuances, thereby ensuring more accurate and robust analyses. Another opportunity lies in researching methods to reduce ”hallucinations” and biases, for instance, through fine-tuning models or adopting techniques for validation and filtering of responses. Moreover, the development of interfaces that make these models more accessible to researchers who are not AI specialists is another area worth exploring. Finally, creating a standardized and robust evaluation metric capable of capturing the complexity of qualitative analysis performed by LLMs represents a promising line of investigation.

### v2

@@ Abstract
The presence of Artificial Intelligence (AI) has been intensifying in supporting various human activities. Generative AI, particularly Large Language Models (LLMs), has demonstrated significant potential for automating complex processes by processing and generating human-like text [1],[2],[3]. These models are trained on extensive textual data and designed to identify complex patterns in natural language, making them versatile tools for tasks such as machine translation, summarization, text editing, chatbot assistance, and even software code generation [4],[5].
One of the most promising applications of LLMs is in qualitative analysis, a research approach focused on exploring and interpreting data to uncover patterns, themes, and categories. Traditional qualitative analysis methods, such as grounded theory or thematic analysis, often require intensive human effort and are prone to subjectivity [6],[7]. Moreover, these methods struggle with scalability, particularly when analyzing large datasets, such as hundreds of interviews or millions of social media posts. LLMs offer transformative solutions to these challenges, automating processes like open coding and theme extraction while providing consistent and reproducible outputs [8],[9].
The integration of LLMs into qualitative research presents transformative opportunities and challenges, particularly in domains such as Software Engineering (SE). As highlighted by Bano et al. [10], LLMs have the potential to address longstanding challenges in qualitative research, such as the time-consuming nature of manual analysis, inconsistencies in interpretation by multiple simultaneous researchers, and scalability limitations. While LLMs can assist in automating tasks such as coding, theme identification, and pattern recognition, they also raise critical concerns related to contextual understanding, reproducibility, and ethical implications. These insights underscore the urgent need for systematic investigations into the use of LLMs in qualitative research.
@@
This paper is structured as follows: Section II provides a brief background on qualitative analysis and LLMs. Section III describes the systematic mapping protocol and results. Section IV presents a summary of contributions and research opportunities. Section V discusses the threats to validity. Section VI presents the main conclusions. Finally, Section VII presents the take away lessons.
@@ II-B Capabilities of LLMs in Overcoming Limitations
LLMs offer transformative solutions to many challenges faced by traditional qualitative research methods. LLMs, such as GPT-4, LLaMA, and ChatGPT, are generative AI systems trained on vast datasets, enabling them to process and generate human-like text [4],[5]. By leveraging billions of parameters, these models excel in identifying patterns, performing thematic analysis, and automating coding processes.
@@
Reducing Subjectivity: While human coding is prone to inconsistency, LLMs provide consistent and reproducible outputs, especially when fine-tuned or used with standardized prompts [8]. However, their adaptability and evolution over time may impact reproducibility. This minimizes bias and ensures greater reliability in qualitative findings, though care must be taken to account for potential variability in outputs across different versions of the models.
@@ II-C Illustrative Examples
Healthcare Case Study: In a study analyzing patient feedback, an open-source LLM was fine-tuned to perform inductive thematic analysis, identifying critical concerns like delays in care and communication breakdowns [9]. This process significantly reduced the time required for analysis while maintainiong accuracy.
@@ II-D Motivation for This Study
Despite their potential, LLMs face limitations that must be addressed for optimal application in qualitative research. For example, they are highly dependent on well-structured prompts, which, if poorly designed, can lead to inaccurate or incomplete outputs [13],[9]. Additionally, LLMs can generate “hallucinations”—fabricated responses that lack basis in the data [14]. These challenges highlight the need for systematic studies to explore best practices, identify gaps, and refine methodologies for integrating LLMs into qualitative research.
@@
The study Leça et al.[15] also converges in this direction by presenting a systematic mapping to investigate how LLMs are typically used in qualitative analysis and how they can be applied in Software Engineering research. The study Leça et al.[15] and this study complement each other in several aspects. However, while in Leça et al.[15] explores benefits (e.g., support in theme and pattern identification and flexibility for researchers) and limitations (e.g., ethical concerns and privacy), this study focuses specifically on the effectiveness of LLMs and the evaluation methods used, including specific metrics and criteria. Together, these studies provide a comprehensive perspective: systematically mapping the state of the art as a guide for future work and offering an in-depth analysis of the contextual implications and challenges in using these technologies. Both contribute to the understanding and applicability of LLMs in qualitative analysis.
@@ III-B Search Strategy
A generic search string was developed using the keywords “LLM” and “Qualitative Analysis”. These keywords were connected using the logical operator AND, while their synonyms were linked with the OR operator. The terms in the search string were selected to ensure broader coverage of relevant studies. The string was tested in various configurations on the Scopus database, and after a calibration process, the final version defined was:
(“LLM” OR “LLMs” OR “Large Language Model” OR “Large Language Models”) AND (“Qualitative Analysis” OR “Qualitative Research” OR “Grounded Theory”)
In addition to Scopus222http://www.scopus.com, searches using the search string were also conducted in the ACM Digital Library333https://dl.acm.org/, IEEExplore444https://ieeexplore.ieee.org/, Web of Science555http://www.isiknowledge.com, and SBC Open Lib databases666https://sol.sbc.org.br/busca/. The searches were carried out during the months of September and October 2024, using filters applied to titles, abstracts, and keywords. No restrictions were placed on the starting publication year. In total, 354 studies were retrieved.
Considering the emerging nature of the research topic and the likelihood of relevant studies not yet undergoing peer review, the Arxiv777https://arxiv.org/ database was also explored using keywords directly, without applying the search string. Arxiv papers were carefully considered due to their potential to provide early access to significant studies that are not yet indexed in major digital libraries like Scopus. Including these papers ensured a comprehensive exploration of the state of the art in LLMs for qualitative analysis.
@@ III-C Selection Criteria
It is important to note that EC1 is presented as the direct inversion of IC1 to emphasize the mutual exclusivity of these criteria, ensuring clarity in identifying studies that analyze the application of LLMs in qualitative analysis. The remaining exclusion criteria (EC2, EC3 and EC4) were defined independently to address specific attributes of the studies, such as language, accessibility, or type of publication.
@@ III-E Conducting
Subsequently, a refinement process was conducted to select the final set of studies. First, duplicate studies were removed, resulting in 253 studies. Next, inclusion and exclusion criteria were applied based on the reading of titles and abstracts, reducing the total to 34 studies. In the following stage, only studies available for download were included, which further reduced the count to 21. The study [20] was a special case. It was retrieved but was excluded due to lack of full-text access. However, in January 2025, the authors published the complete version on ArXiv. We opted to manually include the study to complement the mapping.
If a study did not fully meet the response for criterion DE1 (i.e., the response was “no” or “partially”), the remaining criteria were not evaluated, and the study was excluded. Additionally, if a study did not meet criterion DE11 (“no”), it was excluded for not describing the prompt engineering used, which is essential to ensure reproducibility.
At the end of this process, 8 studies remained, for which comprehensive data extraction was conducted. Table III presents all the studies that were included in this systematic mapping.
@@
| S8 | CollabCoder: A Lower-barrier, Rigorous Workflow for Inductive Collaborative Qualitative Analysis with Large Language Models | [20] |
@@ III-F Results
The analysis revealed that 75% of the included studies (6 out of 8) were published in 2024, with the remaining two in 2023. This reflects the recent emergence of LLMs in qualitative analysis and the growing interest of the scientific community in this field.
Regarding publication venues, most studies were concentrated in healthcare, education, and cultural studies, with no works identified in Software Engineering (SE). This gap is notable given the potential of LLMs to assist in SE tasks such as requirements engineering and user feedback analysis. A possible reason for this absence is the early-stage adoption of LLMs in SE. Future research could explore their integration in SE-specific venues, assessing their impact on traditionally manual processes.
@@
Studies on the use of LLMs for qualitative analysis are concentrated in fields such as healthcare [8],[9], education [14],[17],[13], cultural studies [19], and technological applications [18].
Regarding the primary objectives, three studies evaluated the potential of LLMs to automate qualitative analysis [14],[17],[13]. The study [20] proposed CollabCoder, a workflow integrating LLMs into collaborative qualitative analysis, assisting in code suggestion, discussions, and codebook development. Other studies focused on fully automating qualitative data analysis [8],[19], while [18],[9] compared LLM-assisted qualitative analysis with traditional methods.
@@
The studies examined various LLM models, including ChatGPT [8],[18],[13],[20], LLaMA-2 [9], ATLAS.TI [14], BERT [17] and Sabiá-2 medium [19]. About 50% of the studies made adjustments or adaptations to the models for analysis [14],[8],[17],[9]. However, study [20] used ChatGPT without fine-tuning, relying solely on prompt engineering to assist in qualitative coding.
Regarding data sources, most studies analyzed interviews [14],[9],[13], followed by documents [8],[17], including [20], which used book reviews as qualitative textual data. Other studies examined social media (app user reviews) [18] and song lyrics [19].
@@
All included studies applied techniques and/or methodologies for qualitative analysis, also detailing the prompt engineering used. Theme extraction was combined with content analysis in two studies [14],[18], while others integrated open coding and theme extraction [9],[13]. Two studies combined categorization and theme extraction, with one adopting the Grounded Theory approach [8] and the other using content analysis [17]. Study [20] followed an inductive approach, structuring open coding and theme extraction within a collaborative workflow based on Grounded Theory and Thematic Analysis. Study [19] tested three techniques: categorization, theme extraction, and topic extraction.
LLM application methods varied among studies. Three employed prompt-based instruction [18],[9],[13]. Study [19] combined prompt engineering with automated extraction using BERT, while [17] exclusively used automated extraction with BERT. In [20], ChatGPT provided automated code suggestions, facilitated coder discussions, and assisted in developing final code groupings within an AI-assisted workflow. Other studies applied automated extraction via ATLAS.TI, linked to specific techniques [14], or conducted model training [8].
@@
All seven included studies compared LLM-assisted qualitative analysis with traditional methods. In [14],[8],[17],[9],[13], results were considered equivalent to traditional methods. Study [19] reported superior performance for LLMs, whereas [18] found LLM-assisted analysis to be less effective, particularly compared to human analysis. Study [20] assessed effectiveness through a user study comparing CollabCoder with Atlas.ti Web, revealing that CollabCoder improved efficiency by streamlining coder discussions and decision-making while maintaining high agreement rates. Participants also reported a lower learning curve and better workflow support compared to Atlas.ti Web.
The metrics used to evaluate performance included comparisons between LLM-assisted analysis and human analysis [14],[18],[9],[13], comparisons with human analysis and accuracy [17], and comparisons involving human analysis, accuracy, F1 Score, and precision [8],[19]. In [20], the evaluation included metrics such as Cohen’s Kappa and Agreement Rate to assess coder consensus, demonstrating that the system facilitated alignment and reduced disagreement in the coding process.
@@
All studies identified specific limitations in using LLMs for qualitative analysis. One limitation mentioned was the reliance on well-structured prompts for accurate results [9],[13]. Another recurring limitation was the tendency of LLMs to generate “hallucinations”, producing responses without clear justification from the data [14],[13]. Additionally, inherent model biases were noted, particularly when dealing with sensitive information [9]. Other limitations included difficulties in assigning topics to subjective expressions [19], lack of context sensitivity and emotional nuance [18], and restrictions in interpreting complex meanings [13]. Study [8] also highlighted excessive response over-segmentation and over-categorization. In [20], challenges were reported in achieving coder consensus when dealing with ambiguous data, requiring human intervention for nuanced interpretations.
Further limitations included difficulties in broader and more detailed categorization, leading to a loss of specific details important by human researchers [14], as well as inconsistencies in zero-shot classifications [19]. Another concern was the “echo chamber” effect, with LLMs reproducing pre-trained patterns [18]. Study [20] also highlighted challenges in maintaining coder independence while fostering collaboration, along with risks of premature influence from LLM-generated coding suggestions that could bias human coders.
Regarding recommendations for LLM usage, studies [14],[9],[13] propose that LLMs serve as aids rather than replacements for human analysts, particularly in categorization, with researchers retaining final interpretative authority. Other works emphasize the importance of human collaboration and LLMs role in enhancing accuracy and validity [18],[13]. Study [20] suggests leveraging LLMs to structure and refine codebooks while ensuring human coders make final decisions. It also stresses the need for greater user control over AI-generated suggestions to prevent over-reliance on automated coding.
For future research, there is broad agreement on the need to refine prompt engineering techniques and explore more advanced architectures to enhance LLMs’ textual understanding [17],[18],[19],[13]. Increasing training data volume is also recommended to improve accuracy [17], along with fine-tuning pre-trained models for better topic assignments [19]. Study [20] suggests optimizing prompt design to enhance AI-assisted code generation, refining interaction models for better coder collaboration and include integrating multimodal. Additional research directions include enhancing LLM capabilities to capture semantic and subjective nuances [14]. Study [9] suggestes that expanding open-source models for cost reduction and accessibility, and developing more user-friendly interfaces for researchers.
@@ IV Summary of Contributions and Research Opportunities
This section summarizes the main findings and contributions of this study, as well as the research opportunities identified. The key contributions include:
Mapping of the field:This study provides an overview of the use of LLMs in qualitative analysis, covering various aspects such as: (i) application contexts, (ii) LLM models and data sources, (iii) techniques and methodologies adopted, (iv) metrics used to evaluate effectiveness, and (v) limitations and research opportunities in the current state of the art.
This mapping focused on studies that reported LLM applications for qualitative analysis, detailing the prompt engineering employed—an essential factor for ensuring reproducibility. The findings highlight the recent nature of this field, reflected in the relatively low number of included studies, most of which were published in 2023 and 2024.
However, several research opportunities emerged. While LLMs exhibit performance comparable to or even superior to traditional methods in some cases, improvements are still needed. A key issue is their reliance on well-structured prompts, necessitating further exploration of approaches that enhance LLMs’ textual understanding, allowing for greater flexibility in qualitative analyses.
Additionally, opportunities exist to investigate new architectures and techniques that improve the capture of semantic and subjective nuances, ensuring more precise and robust analyses. Another promising direction involves methods to mitigate “hallucinations” and biases, such as fine-tuning models or employing validation and filtering techniques. The development of user-friendly interfaces for researchers without AI expertise is also a relevant area of exploration. Finally, establishing a standardized and robust evaluation metric capable of capturing the complexity of qualitative analysis performed by LLMs represents a significant avenue for future research.